Can you help me solve the equations ax + by = m and cx + dy = n. somehow the answer is supposed to be those two equations combined into one.
I need an answer asap
everything I have done is wrong and I need to know how to figure them out.

Answer by Alan3354(69443)

Can you help me solve the equations ax + by = m and cx + dy = n.
-----------
Use determinants
|a b -m|
|c d -n|
---------
Det = ab - cd
---
x*Det = b*-n - d*-m = dm - bn
-y*Det = a*-n - c*-m = cm - an
--------
x = (dm - bn)/(ab - cd)
y = (an - cm)/(ab - cd)
===========================
Without determinants:
ax + by = m Eqn 1
cx + dy = n Eqn 2
------
acx + bcy = cm Eqn 1 times c
acx + ady = an Eqn 2 times a
------------------------ Subtract
y*(bc-ad) = cm - an
y = (an - cm)/(ad - bc) (I flipped both NUM and DEN)
========================
ax + by = m Eqn 1
cx + dy = n Eqn 2
------
adx + bdy = dm Eqn 1 times d
bcx + bdy = bn Eqn 2 times b
------------------------ Subtract
x*(ad - bc) = dm - bn
x = (dm - bn)/(ad - bc)